OpenStudy (anonymous):

The 1st electron in the H2 atom is ejected from the atom after absorbing a total amt of Energy = 20.4eV . Calculate its de Broglie wavelenght in A^0 (mass of electron is 9.1 *10^-31 kg)

OpenStudy (anonymous):

@ghazi

OpenStudy (anonymous):

@ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi @ghazi

OpenStudy (ghazi):

ok..wait a sec dude :)

OpenStudy (ghazi):

first of all angstrom= 10^-10 m is that okay?

OpenStudy (ghazi):

then \[E= \frac{ h*c }{ \lambda }\] but it would be great if you can convert this eV

OpenStudy (ghazi):

1eV= 1.60217646 × 10^-19 joules

OpenStudy (anonymous):

but that does nt work..))

OpenStudy (anonymous):

i did..this 1 st time....and let me recheck it..

OpenStudy (anonymous):

mass is given in the question (electron)

OpenStudy (anonymous):

i think we shuld use it...

OpenStudy (ghazi):

\[32.6808* 10^{-19}=\frac{ 6.626*10^{-34}*3*10^8 }{ \lambda }\]

OpenStudy (anonymous):

u would get 60.... some thing

OpenStudy (anonymous):

but the answer shuld be 4.7 A

OpenStudy (anonymous):